If 7 ballerinas dance for 8 hours and lose a total of 20 pounds,
how many more ballerinas dancing would it take to lose a total
of 20 pounds in only 4 hours if the new dancers lost weight only
half as fast as the original 7 ?

Answer by stanbon(75887)

If 7 ballerinas dance for 8 hours and lose a total of 20 pounds,
Rate at which they lost weight was 20/8 = 5/2 = 2.5 lbs per hour or 7
2.5/7 = 0.3571 lbs per hr. per ballerina
---------------------------------------------------------
Note: rate = total loss/[total hours*# of ballarina)
=========================================================
How many more ballerinas dancing would it take to lose a total
of 20 pounds in only 4 hours if the new dancers lost weight only
half as fast as the original 7 ?
---
Let Let "x" be the total # of ballarinas required.
Equations:
(0.3571/2) = 20/[4x]
0.1786 = 5/x
x = 5/0.1786
x = 28 ballarinas would be required
================================
Cheers,
Stan H.